OpenStudy (anonymous):

I need help with 3 questions will someone help me plz.

OpenStudy (anonymous):

Choose the correct slope of the line that passes through the points (-4, 8) and (-3, -6). A. 7 B. -14 C. 0 D. 14

OpenStudy (mertsj):

The slope formula is: \[m=\frac{y _{2}-y _{1}}{x _{2}-x _{1}}\]

OpenStudy (mertsj):

In your problem : \[y _{2}=-6, y _{1}=8, x _{2}=-3, x _{1}=-4\]

OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (-1, 6) and has a slope of -3. A. y - 6 = -3x - 3 B. y - 6 = -3(x + 1) C. y = -3x + 3 D. 3x + y = 3

OpenStudy (mertsj):

Point slope form is: \[y-y _{1}=m(x-x _{1})\]

OpenStudy (anonymous):

So it would be B

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

and the first in is D

OpenStudy (mertsj):

Better do that again.

OpenStudy (anonymous):

B

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

Which of the following will give you the incorrect slope? A. the quantity y subscript two minus y subscript one over the quantity x subscript two minus x subscript one B. the quantity y subscript two minus y subscript one over the quantity x subscript one minus x subscript two C. the quantity y subscript one minus y subscript two over the quantity x subscript one minus x subscript two D. rise over run

OpenStudy (anonymous):

Is it A